# Dimension-independent convergence rates of randomized nets using median-of-means

### v1

@@ 5 Tractability under more general randomization
From [8, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [8, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
However, this refinement offers limited practical improvement: Lemma 6 suggests $|K_{u}(T)|$ grows exponentially in $T$, so the margin $T_{u,m}-(\alpha+\lambda)(m-\tilde{t}_{u})$ is at best $O(\alpha+\lambda)$. Consequently, Consequently, the error bound improves by only a factor of $2^{O(\alpha+\lambda)}$, leaving the asymptotic convergence rate unchanged.
@@
$\leqslant$ $2^{(T-T^{\prime})/\beta}\Big{(}\max\big{(}(T-T^{\prime})/\beta,1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|}+2^{(T-\theta^{\prime}T^{\prime})/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}},$
where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are given by equation (21), $\beta=\alpha+\lambda-d$ and
$D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}}.$
@@
$\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
@@
$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha,\forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
@@
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
@@
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (36)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
@@
where the last step follows from a calculation similar to equation (27).
@@
$t^{T}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
for $t=2^{-1/(\alpha+\lambda-d)}$. To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$. For $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$. Because for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
@@
the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition. By summing this bound over all elements of $\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$, we get
$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}$
@@
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
@@
where we have again applied equation (26). Our conclusion finally follows from equation (36), $t=2^{-1/(\alpha+\lambda-d)}$ and
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
$=$ $\sum_{w\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ w=\{j\mid\alpha^{-}_{j}=d\}\end{subarray}}\Big{(}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
$=$ $\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|}.\qed$
@@
$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (38)

### v2

@@ 5 Tractability under more general randomization
From [9, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [9, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
However, this refinement offers limited practical improvement: Lemma 6 suggests $|K_{u}(T)|$ grows exponentially in $T$, so the margin $T_{u,m}-(\alpha+\lambda)(m-\tilde{t}_{u})$ is at best $O(\alpha+\lambda)$. Consequently, the error bound improves by only a factor of $2^{O(\alpha+\lambda)}$, leaving the asymptotic convergence rate unchanged.
@@
$\leqslant$ $2^{(T-T^{\prime})/\beta}\Big{(}A_{\alpha,\lambda}\max\big{(}(T-T^{\prime})/\beta,1\big{)}+B_{\alpha,\lambda}\Big{)}^{|u|}+2^{(T-\theta^{\prime}T^{\prime})/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}},$
where $\beta=\alpha+\lambda-d$, $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are given by equation (21), and
$D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}},$ (35)
with the sum from $1$ to $\alpha-d$ interpreted as 00 when $\alpha=d$.
@@
$\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
@@
$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
@@
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
@@
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
@@
where the last step follows from a calculation similar to equation (25) and (27).
@@
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant t^{-T}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (38)
for $t=2^{-1/(\alpha+\lambda-d)}$. To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$. For $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$. Because for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
@@
the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition. This implies $|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\times|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ and
$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}},$
@@
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
@@
where we have applied equation (26) once more. Our conclusion finally follows from equation (37), $t=2^{-1/(\alpha+\lambda-d)}$ and
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
$=$ $\sum_{w\subseteq u}\Big{(}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ \alpha_{j}=d\ \forall j\in w\\ \alpha^{-}_{j}>0\ \forall j\in u\setminus w\end{subarray}}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (35). ∎
@@
$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (39)